OpenStudy (anonymous):

Find a quadratic function to model the values in the table. Predict the value of y for x = 6 x= 1,0,3 y= 2,2,-10

OpenStudy (goformit100):

y^2 - 36 = 0 is the required equation.

OpenStudy (anonymous):

the answer choices are A. y = -2x2 + 2x -2; –58 B. y = 2x2 – 2x -2; 60 C. y = 2x2 – 2x -2; 58 D. y = -2x2 + 2x +2; –58

OpenStudy (goformit100):

Shaw your working

OpenStudy (anonymous):

i dont know how to solve the problem, that's why im showing the choices. i need a solution please?